OpenStudy (anonymous):

1 + cos(2x) = cot(x)sin(2x)

OpenStudy (anonymous):

LHS= 1+COS2X=1+2COS^2X -1=2COS^2X RHS= COTX.SIN2X=COSX/SINX.2SINXCOSX=2COS^2X LHS=RHS...........